# Intelligent Collaborative Optimization for Rubber Tyre Film Production Based on Multi-path Differentiated Clipping Proximal Policy Optimization

### v1

@@ I Introduction
Figure 1: Calendering Equipment in the Rubber Tyre Production Line.
@@
Digital twins facilitate real-time mapping of physical systems, supporting parameter optimization and anomaly diagnosis, while DRL techniques, including Proximal Policy Optimization (PPO)[6] and Twin Delayed Deep Deterministic Policy Gradient (TD3)[7], have shown strong performance in continuous control and multi-agent collaboration. Additionally, Transformer and graph neural networks (GNNs) enhance high-dimensional state representation and cross-domain generalization. Nevertheless, tyre production remains a challenging domain due to its high dimensionality, strong coupling, and nonlinear dynamics, demanding more robust and adaptive collaborative algorithms.
@@ II-A Deep Learning for Industrial Prediction
In industrial image processing, Convolutional Neural Networks (CNNs)[8] are widely applied in tasks such as surface defect detection and target positioning. Evolving from LeNet[9] to AlexNet[10], VGG[11], and ResNet[12], these architectures have significantly advanced feature extraction for visual inspection. Lightweight networks like MobileNet[13] and SqueezeNet[14] enable real-time edge applications, though with accuracy-efficiency trade-offs. Recently, Vision Transformers (ViT)[15] have demonstrated remarkable capabilities in capturing long-range dependencies, despite higher computational costs.
@@ II-B Reinforcement Learning for Industrial Control
Hierarchical RL (HRL)[31] and meta-RL[32] offer structured task decomposition and fast adaptation to new industrial scenarios, supporting multi-scale decision-making and reducing retraining needs.
@@ IV-A Data Acquisition
To thoroughly investigate the relationship between rubber film width and various machine/environmental parameters, authentic process data were collected from the inner-liner production line. The MESNAC [^1] https://en.mesnac.com/, a key strategic partner and major supplier of intelligent manufacturing systems to the global tyre industry, provided continuous in-situ measurements via industrial IoT sensors and monitoring devices. The company’s strong R&D capabilities ensured the reliability and precision of the dataset used in this study. As illustrated in Fig. 1, the manufacturing process begins with compound plasticization in an extruder. The material is then discharged onto a conveyor and fed into a calender to form a continuous sheet. This sheet is subsequently cropped to the required width by edge-trimming knives, drawn through pull rolls, and finally delivered to the take-off conveyor. Thermal expansion and contraction cause the width and thickness of the trimmed and drawn strip to fluctuate until the sheet stabilizes at the take-off section. A critical constraint arises from the equipment setup: width and thickness gauges are installed several meters downstream on the take-off belt, introducing a measurement delay. This delay not only prevents real-time monitoring of the current process state but also leads to material waste during product-specification changes. Consequently, the continuous process data collected from this line (Fig. 3) provide a valuable foundation for developing intelligent predictive models. According to plant engineers, film width and thickness are governed by the machine parameters listed in Tables I and II. By tuning these parameters, it is possible to achieve concurrent optimization of both width and thickness.
@@ IV-B Experimental Results Of Prediction
Width 1.681 2.32 0.547
Thickness 0.156 0.219 0.456
To systematically evaluate the effectiveness and generalization of the proposed prediction model, this study conducts an empirical comparative analysis using tyre film width prediction as a representative production task. All models were trained under identical hyperparameters, including a learning rate of 0.001 and a batch size of 1024, and for 100 epochs per run, with results averaged over 10 independent trials. Four models were used for comparison, including BILSTM[34], Linear Regression(LR), TCN[35] and Informer[36]. The models were evaluated on four metrics: MAE, RMSE, Error Variance(EV, the variance of prediction errors), and Qualification Rate (QR). The comparative experimental results are summarized in Table IV.
@@ IV-C Experimental Results Of Optimization
For width control, a looser clipping bound allows for larger-magnitude adjustments, whereas for thickness, a tighter bound prevents aggressive updates to ensure stable fine-tuning. This study investigates the impact of training steps per episode on the optimization performance of the MPD-PPO agent. We systematically compared configurations of 50 and 100 training steps per episode to evaluate their effects on convergence behavior and solution quality, where at each time step the agent selects an action. The experimental outcomes after 50 and 100 training episodes are presented in Fig. 4, Fig. 5, and Table V.
@@
Train Step Target (Width, Thickness) ($mm$) Optimize Step
50 (480, 3.0) 27
@@
At 50 training steps, the optimization results demonstrate considerable effectiveness, though minor oscillations are observed in localized regions. By 100 training steps, the agent exhibits significantly improved performance, consistently achieving optimization targets within a maximum of 20 steps across all evaluated episodes. These results indicate robust convergence and high sample efficiency of the proposed method.
@@ IV-D1 The Influence Of Multi-Branch Network
In this experiment, we replaced the original single-network architecture of the PPO algorithm with a two-branch network, comprising separate width and thickness action networks to investigate the efficacy of the branched architecture. Representative results for the typical operating condition of 480 mm width and 3.0 mm thickness are presented for comparative analysis. Table VI summarizes the quantitative results. Figures 6 and 7 present the performance of PPO with Multi-Branch Network and standard PPO, respectively, following the MPD-PPO results in Figure 5.
@@
TABLE VI: Optimization Results Of Remove Differentiated Clip.
Algorithm Optimize Step
MPD-PPO 16
@@ IV-D2 The influence of Differentiated Clipping
This experiment sets identical clipping thresholds for both branches to investigate the efficacy of the differentiated clipping strategy. Representative results for a width of 480 mm and a thickness of 3.0 mm are presented for comparative analysis. The corresponding quantitative results are summarized in Table VII. Compared to the standard MPD-PPO in Fig. 5, Fig. 8 displays the MPD-PPO No Differentiated Clipping results.
@@
Algorithm Optimize Step
MPD-PPO 16
MPD-PPO No Differentiated Clipping 87
@@ IV-D3 Impact of Reward-Function Modules on Optimization Performance
Algorithm Optimize Step
@@
Reward 3 (Excludes The Steady-State Reward) 21
Reward 4 (The Complete Reward Function) 15
@@ IV-E Experimental Validation on a Real-World Rubber Tyre Production Line
To rigorously validate the algorithm’s effectiveness under genuine industrial conditions, this study conducted testing in collaboration with MESNAC, on one of its operational production lines. Process data were collected in real-time via PLC communication at a frequency of 0.2 ms intervals, directly from sensors and control systems. This in-situ validation strategy guarantees a high degree of fidelity to the actual production environments, thereby providing dependable evidence for assessing the algorithm’s performance. The experimental results are presented in Fig. 10 and Fig. 11. Experimental results from the production line demonstrate the robustness of the agent’s control strategy. In response to both positive and negative setpoint adjustments in width, the agent generated appropriate control actions that effectively guided the system state toward the target. The regulation process exhibited no significant overshoot, indicating favorable convergence characteristics.
@@ V Conclusion
This study presented a novel intelligent control framework that seamlessly integrates deep reinforcement learning with predictive modeling to address multi-objective parameter optimization in complex industrial processes.

### v2

@@ I Introduction
Figure 1: Rubber Tyre Film Manufacturing Equipment.
@@
Digital twins facilitate real-time mapping of physical systems, supporting parameter optimization and anomaly diagnosis, while DRL techniques, including Proximal Policy Optimization (PPO) [6] and Twin Delayed Deep Deterministic Policy Gradient (TD3) [7], have shown strong performance in continuous control and collaborative optimization. Additionally, Transformer and graph neural networks (GNNs) enhance high-dimensional state representation and cross-domain generalization. Nevertheless, tyre production remains a challenging domain due to its high dimensionality, strong coupling, and nonlinear dynamics, demanding more robust and adaptive collaborative algorithms.
@@ II-A Deep Learning for Industrial Prediction
In industrial image processing, Convolutional Neural Networks (CNNs) [8] are widely applied in tasks such as surface defect detection and target positioning. Evolving from LeNet [9] to AlexNet [10], ResNet [11], and DenseNet [12], these architectures have significantly advanced feature extraction for visual inspection. Lightweight networks like MobileNet [13] and ShuffleNet [14] enable real-time edge applications, though with accuracy-efficiency trade-offs. Recently, Vision Transformers (ViT) [15] have demonstrated remarkable capabilities in capturing long-range dependencies, despite higher computational costs.
@@ II-B Reinforcement Learning for Industrial Control
Model-based RL (MBRL) [31] and Offline RL [32] offer improved sample efficiency through internal simulation and effective utilization of historical datasets, supporting multi-scale decision-making and reducing retraining needs.
@@ III-D4 Steady-State Reward
The reward coefficients were calibrated through sensitivity analysis to ensure balanced collaborative optimization, with total rewards clipped to a predefined range for numerical stability during training. This structured reward design effectively integrates tracking precision, behavioral smoothness, and steady-state performance, addressing critical requirements of industrial control systems.
@@ IV-A Data Acquisition
To thoroughly investigate the relationship between rubber film width and various machine/environmental parameters, real-world process data were collected from the inner-liner production line. MESNAC [^1] https://en.mesnac.com/, a key strategic partner and major supplier of intelligent manufacturing systems to the global tyre industry, provided continuous in-situ measurements via industrial IoT sensors and monitoring devices. The company’s strong R&D capabilities ensured the reliability and precision of the dataset used in this study. As illustrated in Fig. 1, the manufacturing process begins with compound plasticization in an extruder. The material is then discharged onto a conveyor and fed into a calender to form a continuous sheet. This sheet is subsequently cropped to the required width by edge-trimming knives, drawn through pull rolls, and finally delivered to the take-off conveyor. Thermal expansion and contraction cause the width and thickness of the trimmed and drawn strip to fluctuate until the sheet stabilizes at the take-off section. A critical constraint arises from the equipment setup: width and thickness gauges are installed several meters downstream on the take-off belt, introducing a measurement delay. This delay not only prevents real-time monitoring of the current process state but also leads to material waste during product-specification changes. Consequently, the continuous process data collected from this line (Fig. 3) provide a valuable foundation for developing intelligent predictive models. According to plant engineers, film width and thickness are governed by the machine parameters listed in Tables I and II. By tuning these parameters, it is possible to achieve concurrent optimization of both width and thickness.
@@ IV-B Experimental Results Of Prediction
[…]
To systematically assess the effectiveness and generalization of the proposed prediction model, this study conducts an empirical comparative analysis using tyre film width prediction as a representative production task. All models were trained under identical hyperparameters, including a learning rate of 0.001 and a batch size of 1, 024, and for 100 epochs per run, with results averaged over 10 independent trials. Four models were used for comparison, including BILSTM[34], Linear Regression (LR), TCN[19] and Informer[35]. The models were evaluated on four metrics: MAE, RMSE, and Qualification Rate (QR). The comparative experimental results are summarized in Table IV.
@@ IV-C Experimental Results Of Optimization
For width control, a looser clipping bound allows for larger-magnitude adjustments, whereas for thickness, a tighter bound prevents aggressive updates to ensure stable fine-tuning. This study investigates the impact of training steps per episode on the optimization performance of the MPD-PPO agent. We systematically compared configurations of 50 and 100 training steps per episode to evaluate their effects on convergence behavior and solution quality, where at each time step the agent selects an action. To enhance the robustness of the findings, we conducted five independent experimental trials and consolidated the resulting optimization curves into an intuitive interval representation for visualization. The experimental outcomes after 50 and 100 training episodes are presented in Fig. 4, Fig. 5, and Table V. It is noteworthy that the thickness curve appears wider post-convergence in the figure. This is a visual effect resulting from the high resolution of the y-axis scale, which is necessitated by its small absolute numerical range, rather than an indication of increased actual volatility.
@@
Train Step Target (Width, Thickness) ($mm$) Average Optimize Step
50 (480, 3.0) 23
@@
At 50 training steps, the optimization results demonstrate considerable effectiveness, though minor oscillations are observed in localized regions. By 100 training steps, the agent exhibits significantly improved performance, consistently achieving optimization targets within approximately 20 steps across all evaluated episodes. These results indicate robust convergence and high sample efficiency of the proposed method.
@@ IV-D1 The Influence Of Multi-Branch Network
In this experiment, we replaced the original single-network architecture of the PPO algorithm with a two-branch network, comprising separate width and thickness action networks to investigate the efficacy of the branched architecture. Representative results for the typical operating condition of 480 mm width and 3.0 mm thickness are presented for comparative analysis. Table VI summarizes the quantitative results. Figures 6 and 7 present the performance of PPO with Multi-Branch Network and standard PPO, respectively, following the MPD-PPO results in Fig 5.
@@
TABLE VI: Optimization Results Of Removing Differentiated Clip.
Algorithm Average Optimize Step
MPD-PPO 22
@@ IV-D2 The influence of Differentiated Clipping
This experiment sets identical clipping thresholds for both branches to investigate the efficacy of the differentiated clipping strategy. Representative results for a width of 480 mm and a thickness of 3.0 mm are presented for comparative analysis. The corresponding quantitative results are summarized in Table VII. Compared to the standard MPD-PPO in Fig. 5, Fig. 8 displays the MPD-PPO without Differentiated Clipping results.
@@
Algorithm Average Optimize Step
MPD-PPO 22
MPD-PPO Without Differentiated Clipping 87
@@ IV-D3 Impact of Reward-Function Modules on Optimization Performance
Algorithm Average Optimize Step
@@
Reward 3 (Excludes The Steady-State Reward) 22
Reward 4 (The Complete Reward Function) 16
@@ IV-E Experimental Validation on a Real-World Rubber Tyre Production Line
To rigorously assess the algorithm’s effectiveness under real-world industrial conditions, this study conducted testing in collaboration with MESNAC, on one of its operational production lines, the same industrial equipment as illustrated in Fig. 1. Process data were collected in real-time via PLC communication at a frequency of 0.2 ms intervals, directly from sensors and control systems. This in-situ validation strategy guarantees a high degree of fidelity to the actual production environments, thereby providing dependable evidence for assessing the algorithm’s performance. The experimental results are presented in Fig. 10 and Fig. 11. Experimental results from the production line demonstrate the robustness of the agent’s control strategy. In response to both positive and negative setpoint adjustments in width, the agent generated appropriate control actions that effectively guided the system state toward the target. The regulation process exhibited no significant overshoot, indicating favorable convergence characteristics.
@@ V Conclusion
In this research, we propose a novel intelligent control framework that seamlessly integrates deep reinforcement learning with predictive modeling to address multi-objective parameter optimization in complex industrial processes.